# Cartan–Thullen theorem and Levi problem in context of generalised convexity

### v2

@@ 1.3.1. Families satisfying maximum principle
    there exists a non-negative, proper function $p$ on $\Omega$, bounded on compactae, such that
@@
As we have already noticed in Remark 1.15, when $\mathcal{F}$ is consists of real-parts of a complex algebra, then the conditions (i) of Theorem 1.11, where $\mathcal{A}=\mathcal{F}$, and (i) of Theorem 1.18 are equivalent, see Proposition 4.9. For a characterisation of (i) of Theorem 1.18, see Proposition 4.7.
@@
We shall that the cone $\mathcal{F}$ of functions on a topological space $\Omega$ is local, whenever for any family $(U_{i})_{i\in I}$ of distinct connected components of $\Omega$ and any family $(f_{i})_{i\in I}\subset\mathcal{F}$, there exists $f\in\mathcal{F}$ such that
@@
1.4. Examples
We shall now exhibit several examples how Theorem 1.11 and Theorem 1.18 apply to various settings.
1.4.1. Holomorphic convexity
@@
The following theorem, which is implies Theorem 1.2, follows from Theorem 1.11.
@@
    $\Omega$ is a $\mathcal{B}$-space.
Let us now discuss the relation between the above theorem and Theorem 1.11.
The condition that $\Omega$ is an $\mathcal{B}$-space implies readily that it is a domain of existence, which in turn trivially implies that it is a domain of holomorphy. Similarly, the condition that $\Omega$ is complete with respect to $\mathcal{B}$ shows that it is a domain of holomorphy. This, in turn, is known to imply (i) of Theorem 1.2 by a simple argument involving power series expansion, see [22, Theorem 3.4.5, p. 153].
Therefore Theorem 1.23 implies Theorem 1.2, up to the above mentioned argument. It also implies the modification of (iv) to (iii) of Theorem 1.23.
Let us note, that we may as well employ Theorem 1.18 directly to the setting of an open subset of $\mathbb{C}^{n}$ and the cone of plurisubharmonic functions. Then we recover also [20, Theorem 4.1.19., p. 236].
1.4.2. Stein spaces
@@
    $\Omega$ is a $\mathcal{B}$-domain.
Let us recall, similarly as in the case of open subsets of $\mathbb{C}^{n}$, there exists a characterisation of Stein manifolds that relies on the existence of strictly plurisubharmonic exhaustion function, see [19, Theorem 5.1.6., p. 117].
Let us stress that, up to the best of our knowledge, the conditions (iii), (iv) and (vi) are new. However, we remark that [19, Theorem 5.4.2., p. 138] is similar to the completeness condition and [19, Lemma 5.3.7., p. 133] implies (v). The condition (iii) shows that the intention of Stein, the founder of the notion, to name spaces that satisfy the conditions of Definition 1.24 as holomorphically complete, see [35], was accurate.
1.4.3. Plurisubharmonic functions
When $\mathcal{F}$ is the cone of plurisubharmonic functions, then Theorem 1.18 allows us to prove [20, Theorem 4.1.19., p. 236].
1.4.4. Subharmonic functions
When $\mathcal{F}$ is the cone of subharmonic functions, then Theorem 1.18 allows us to prove [20, Theorem 3.2.31., p. 165].
1.4.5. Convex functions
For a characterisation for the space of convex functions, see [20, Theorem 2.1.25., p. 58] and its corollary [20, Corollary 2.1.26., p. 59], which can again be inferred from Theorem 1.18.
1.5. Gelfand transform and relation to convexity
The notion of convexity that we study here is naturally linked to the classical convexity via the Gelfand transform $\Phi\colon\Omega\to\mathcal{A}^{*}$, see Section 3, that is defined by the formula
@@ 1.5. Gelfand transform and relation to convexity
Using the Gelfand transform, we can transfer linear notions to the setting of the space $\Omega$, in a similar way to transference of the notion of convex hull, see Section 4, Definition 4.4. This allows to reduce studying $\Omega$ and $\mathcal{A}$ to studying subsets of a linear space and linear, continuous functionals on that space.
1.6. Relation to theorems of functional analysis
@@
Let us note the following analogies of Theorem 1.11 and Theorem 1.18 with classical results of functional analysis, see e.g. [31]. These results include the Krein–Šmulian theorem, [23], the Banach–Alaoglu theorem and the Mazur theorem. Indeed, that completeness of $\Omega$ is equivalent to Theorem 1.11, (i), follows from the Banach–Alaoglu theorem, and the Banach–Steinhaus uniform boundedness principle and is an analogue of the Krein–Šmulain theorem, Theorem 5.4. Moreover, that completeness of $\Omega$ implies (i) of Theorem 1.18 follows from the Mazur theorem.
1.7. Applications
Let us mention that generalised convexity has found its applications in a study of the monopolist’s problem by Figalli, Kim and McCann [12] and by McCann and Zhang in [25].
@@ 1.7. Applications
An application of the results obtained in this paper is concerned with a generalisation of martingale transport of measures, which we investigate in [3]. Theorem 1.11 is used there to provide a characterisation of spaces that admit an exhaustion function. In the setting of such spaces in [3] we extend the results of De March and Touzi [9], Obłój and Siorpaes [27], Ghoussoub, Kim, Lim [14] to the setting of martingales with respect to a linear space of functions. Let us mention also [8], where we study a related concept of ordering of measures with respect to a cone of functions.
@@
The aforementioned extension allows for a new type of localisation-type results. Another type of localisation has been studied in [5]. This approach is related to optimal transport of vector measures [7], [6] and extensions of Lipschitz maps [4].
1.8. Structure of article
In Section 2 we provide several definitions and lemmata that will be of use subsequently.
In Section 3 we define the Gelfand transform and show that its basic properties.
In Section 4 we provide a definition of convexity with respect to a convex cone of functions. In Proposition 4.9 show a characterisations of closed convex hulls with respect to a linear space of real-parts of a complex algebra.
In Section 5 we prove the main results, Theorem 1.11, Theorem 1.18 and Theorem 1.23. We also elaborate on the analogy of Theorem 1.11 and the Krein–Šmulian theorem.
2. Preliminaries
@@
We shall say that a topological space $\Omega$ is exhaustible by compactae if there exists a sequence $(K_{i})_{i=1}^{\infty}$ of compact subsets of $\Omega$ such that
@@
It is readily seen that the condition that $\Omega$ is exhaustible by compactae is equivalent to $\Omega$ being $\sigma$-compact and locally compact.
@@
The family of semi-norms $(\lVert\cdot\rVert_{\mathcal{C}(K)})_{K\in\mathcal{K}}$ on $\mathcal{A}$ defines a locally convex topology on $\mathcal{A}$, which we shall denote below by $\tau_{\mathcal{K}}$.
@@
If $\Omega$ is $\sigma$-compact, then $\mathcal{A}$ in $\tau_{\mathcal{K}}$ is metrisable, as the topology is induced by a countable separating family of semi-norms [31, Remark 1.38, c), p. 28]. If moreover $\Omega$ is separable, then for each compact set $K\subset\Omega$ the space $\mathcal{C}(K)$ is separable, as follows by the Stone–Weierstrass theorem, so that $\mathcal{A}$ is separable. Therefore separability and $\sigma$-compactness of $\Omega$ imply that $\mathcal{A}^{*}$ in the weak* topology is metrisable on its compact subsets.
@@
Let $\Omega$ be a set. Let $\mathcal{A}$ be a set of functions on $\Omega$ with values in (,]($\infty$,$\infty$]( - ∞ , ∞ ]. A set $\mathcal{K}$ of functions on $\Omega$ is said to be:
@@
    a convex cone whenever $\alpha f+\beta g\in\mathcal{K}$ for any $f,g\in\mathcal{K}$ and any numbers $\alpha,\beta\geq 0$;
@@
    stable under suprema provided that the supremum $\sup\{f_{i}\mid i\in I\}$ of any family $(f_{i})_{i\in I}\subset\mathcal{K}$ belongs to $\mathcal{K}$;
@@
    separating points of $\Omega$ whenever for any two distinct $\omega_{1},\omega_{2}\in\Omega$, there exist $f,g\in\mathcal{K}$ such that $f(\omega_{1})\neq g(\omega_{2})$.
Let us stress that we allow the functions to take value +$\infty$+ ∞. Note however that this is not allowed for linear subspaces.
@@
(2.1) $f(\omega)=\sup\{a_{i}(\omega)\mid i\in I\}\text{ for all }\omega\in\Omega.$
Proof.
@@
Clearly, $\mathcal{F}\supset\mathcal{G}$. Since $\mathcal{G}$ is a complete lattice cone, $\mathcal{F}=\mathcal{G}$. ∎
@@
If $Y$ is the space all linear, continuous functionals on a linear topological space $X$, then the complete lattice cone generated by $Y$ is the set of all convex, lower semi-continuous functions on $X$. We refer the reader to [29, Chapter 3, p. 13] for a proof of this fact.
@@
Let $\mathcal{F}$ be the complete lattice cone of functions on $\Omega$ that contains constants. Let $p\in\mathcal{F}$. Let $\xi\colon\mathbb{R}\to\mathbb{R}$ convex and non-decreasing. Then the composition $\xi(p)$ belongs to $\mathcal{F}$.
Proof.
@@
For any $j\in J$, there exist $\alpha_{j}\geq 0$ and $\beta_{j}\in\mathbb{R}$ such that $\lambda_{j}(t)=\alpha_{j}t+\beta_{j}$ for all $t\in\mathbb{R}$, so that $\alpha_{j}p+\beta_{j}\in\mathcal{F}$ for each $j\in J$. Thus $\xi(p)\in\mathcal{F}$. ∎
@@
Let $\Omega$ be a topological space. A map $p\colon\Omega\to\mathbb{R}$ is said to be proper whenever preimages $p^{-1}(Z)$ of compact sets in $Z\subset\mathbb{R}$ are compact sets in $\Omega$. Let $\mathcal{F}$ be a cone of functions on $\Omega$. We say that
@@
3. Gelfand transform
Given a topological space $\Omega$ and a linear space $\mathcal{A}$ of continuous functions on $\Omega$, one may consider the space of all linear functionals on $\mathcal{A}$ that are evaluations at points of $\Omega$. Let $\Phi\colon\Omega\to\mathcal{A}^{*}$ be given by the formula $\Phi(\omega)(a)=a(\omega)$ for $\omega\in\Omega$ and $a\in\mathcal{A}$. We shall call $\Phi$ the Gelfand transform, in analogy to the notion studied in the theory of Banach algebras, cf. [38].

### v3

@@ 1.4.1 Cones satisfying the maximum principle
    there exists a non-negative, proper function $p$ on $\Omega$, bounded on compacta, such that
@@
As we have already noted in Remark 1.15, when $\mathcal{F}$ consists of real-parts of a complex algebra, then the conditions (i) of Theorem 1.11, where $\mathcal{A}=\mathcal{F}$, and (i) of Theorem 1.18 are equivalent, see Proposition 4.13. For a characterisation of (i) of Theorem 1.18, see Proposition 4.10.
@@
We shall say that the cone $\mathcal{F}$ of functions on a topological space $\Omega$ is local, whenever for any family $(U_{i})_{i\in I}$ of distinct connected components of $\Omega$ and any family $(f_{i})_{i\in I}\subset\mathcal{F}$, there exists $f\in\mathcal{F}$ such that
@@
1.4.2 Equivalence of the Bremermann–Lelong lemma and the positive resolution of the Levi problem
In the field of complex analysis of several complex variables, the lemma of Bremermann and of Lelong [4, Theorem 2] states that if $\Omega\subset\mathbb{C}^{n}$ is a domain of holomorphy, then any plurisubharmonic function on $\Omega$ is a Hartogs function. The set of Hartogs functions on $\Omega$, cf. [4, Section 2], is the smallest convex cone of functions on $\Omega$ that contains the logarithms of moduli of holomorphic functions on $\Omega$ and is closed under taking:
  1. (i)
    locally bounded suprema,
  2. (ii)
    monotonically decreasing limits.
Moreover, the function belongs to the set of Hartogs functions on $\Omega$ if and only if its restrictions to any open, precompact subset $U$ of $\Omega$ is a Hartogs function on $U$, and any upper semi-continuous regularisation of a Hartogs function is again a Hartogs function.
We shall show in Theorem 6.2 that our methods allow to prove that the positive resolution of the Levi problem is equivalent to the lemma of Bremermann and Lelong.
1.5 Examples
We shall now exhibit several examples of how Theorem 1.11 and Theorem 1.18 apply to various settings.
1.5.1 Holomorphic convexity
@@
The following theorem, which implies Theorem 1.2, follows from Theorem 1.11.
@@
    $\Omega$ is a $(\mathcal{B},d)$-space, where $d$ is any metric on $\Omega$ that generates the standard topology.
Let us now discuss the relationship between the above theorem and Theorem 1.11.
The condition that $\Omega$ is a $(\mathcal{B},d)$-space readily implies that it is a domain of existence, see Example 1.8, which in turn trivially implies that it is a domain of holomorphy. Similarly, the condition that $\Omega$ is complete with respect to $\mathcal{B}$ shows that it is a domain of holomorphy. This is known to imply (i) of Theorem 1.2 by a simple argument involving power series expansion, see [25, Theorem 3.4.5, p. 153]. Let us stress that this argument is not needed to prove the equivalence in Theorem 1.23 above – it is only needed if we want to prove the original version of the Cartan–Thullen theorem. Moreover, the implications in our proof rely on the theorems of functional analysis: the Banach–Alaoglu theorem and the Banach–Steinhaus uniform boundedness theorem. Our proof thus allows to locate the classical Cartan–Thullen theorem in a broader context – as an instance of a general result in generalised convexity.
Moreover, Theorem 1.23 also implies the modification of (iv) to (iii) of Theorem 1.23.
Let us note, that we may as well employ Theorem 1.18 directly to the setting of an open subset of $\mathbb{C}^{n}$ and the cone of plurisubharmonic functions. Then we recover also [23, Theorem 4.1.19., p. 236].
1.5.2 Stein spaces
@@
    $\Omega$ is a $(\mathcal{B},d)$-space for any metric $d$ that generates the same topology as $\tau(\mathcal{B})$.
Let us recall that, similarly as in the case of open subsets of $\mathbb{C}^{n}$, there exists a characterisation of Stein manifolds that relies on the existence of strictly plurisubharmonic exhaustion function, see [22, Theorem 5.1.6., p. 117].
Let us stress that, to the best of our knowledge, the conditions (iii), (iv) and (vi) of Theorem 1.25 are new. However, we remark that [22, Theorem 5.4.2., p. 138] is similar to the completeness condition and [22, Lemma 5.3.7., p. 133] implies (v). The condition (iii) shows that the intention of Stein, the founder of the notion, to name spaces that satisfy the conditions of Definition 1.24 as holomorphically complete, see [40], was accurate.
1.5.3 Plurisubharmonic functions
When $\mathcal{F}$ is the cone of plurisubharmonic functions, then Theorem 1.18 reproves [23, Theorem 4.1.19., p. 236].
1.5.4 Subharmonic functions
When $\mathcal{F}$ is the cone of subharmonic functions, then Theorem 1.18 reproves [23, Theorem 3.2.31., p. 165].
1.5.5 Convex functions
An analogous characterisation for the space of convex functions, see [23, Theorem 2.1.25., p. 58] and its corollary [23, Corollary 2.1.26., p. 59], can again be inferred from Theorem 1.18.
1.6 Gelfand transform and relation to convexity
The notion of convexity that we study here is naturally linked to the classical convexity via the Gelfand transform $\Phi\colon\Omega\to\mathcal{A}^{*}$, which is defined by the formula
@@ 1.6 Gelfand transform and relation to convexity
Using the Gelfand transform, we can transfer linear notions to the setting of the space $\Omega$, in a similar way to transference of the notion of convex hull, see Section 4, Definition 4.6. This allows for a reduction of investigations of an abstract space $\Omega$ equipped with a linear space $\mathcal{A}$ of functions on $\Omega$ to studying subsets of a linear space and linear, continuous functionals on that subset.
1.7 Methods of proofs – functional analytic ingredient
The main novelty in our approach lies in the use of the classical theorems of functional analysis. We equip space the $\mathcal{A}$ with the compact-open topology. This is a topology generated by the family of semi-norms
$\lVert a\rVert_{\mathcal{C}(K)}=\sup\{\lvert a(\omega)\rvert\mid\omega\in K\},$
where $K$ are compact subsets of $\Omega$. Let us assume that $\mathcal{A}$ is closed. The sets
$\{\omega\in\Omega\mid a(\omega)\leq C\sup\lvert a\rvert(K)\},$ (1.1)
considered in (i) of Theorem 1.11 for compact $K\subset\Omega$ and $C\geq 0$, consists precisely of these continuous functionals on $\mathcal{A}$ that are continuous with respect to the semi-norm $\lVert\cdot\rVert_{\mathcal{C}(K)}$, and which are evaluations at the points of $\Omega$. This is to say
$\{\omega\in\Omega\mid\{a(\omega)\leq C\sup\lvert a\rvert(K)\}=\Phi^{-1}(R\cap\Phi(\Omega)),$
where $R$ is a polar set of a neighbourhood of zero in $\mathcal{A}$. The Banach–Alaoglu theorem shows that $R$ is a compact set in $\mathcal{A}^{*}$. Note that $\Phi$ is a homeomorphism onto its image. Moreover, completeness of $\Omega$ with respect to $\mathcal{A}$ is equivalent to the completeness of $\Phi(\Omega)\subset\mathcal{A}^{*}$, and thus also equivalent to its closedness. Thus, we see that completeness implies that the sets of the form (1.1) are compact.
The converse implication is a consequence of the Banach–Steinhaus uniform boundedness principle. Indeed, a Cauchy sequence of elements in $\Omega$ induces the sequence of functionals in $\mathcal{A}^{*}$ – evaluations at the points of the sequence. These functionals are pointwise bounded and therefore there is a neighbourhood of zero of $\mathcal{A}$ on which the sequence is bounded. This means that the original Cauchy sequence belonged to a set of the form (1.1), whose compactness implies convergence of the considered sequence.
In Lemma 5.8 we present an alternative proof of one of the above implications, under an additional assumption that $\mathcal{A}$ consists of real-parts of a complex algebra. The technique we use in this alternative proof is not new – however, it has not been applied previously in the context of generalised convexity.
The equivalence of (i) of Theorem 1.11 and that $\Omega$ is an $(\mathrm{cl}\mathcal{A},d)$-space is an application of the technique usually employed in the setting of complex analysis of several variables. We include the proof for the sake of completeness.
Another novelty in our developments is the proof of the equivalences of (i)-(iv) in Theorem 1.11. As we have mentioned already the previous approaches to the Levi problem used involved analytical or homological tools. In our approach, the desired function is constructed directly in an elementary way that relies on the $\sigma$-compactness assumption on the space $\Omega$. The construction in presented in detail in Lemma 5.1.
The proof of Theorem 1.18 follows similar steps to the proof of the equivalences of (i)-(iv) in Theorem 1.11.
1.7.1 Relation to theorems of functional analysis
@@
Let us note further analogies of Theorem 1.11 and Theorem 1.18 with classical results of functional analysis, the statements of which can be found in [36]. The equivalence of completeness and Theorem 1.11, (i) is an analogue of the Krein–Šmulain theorem, Theorem 5.6. Moreover, that completeness of $\Omega$ implies (i) of Theorem 1.18 follows from the Mazur theorem.
1.8 Applications
@@
An application of the results obtained in this paper is concerned with a generalisation of martingale transport of measures, which we investigate in [11] and in [12]. Theorem 1.11 is used there to provide a characterisation of spaces that admit an exhaustion function, whose existence is an assumption in a generalisation of the Strassen theorem proven in [11]. In the setting of such spaces in [11] we extend the results of De March and Touzi [14], Obłój and Siorpaes [30], Ghoussoub, Kim, Lim [19] to the setting of martingales with respect to a linear space of functions. Let us also mention [10], where we study a related concept of ordering of measures with respect to a cone of functions.
@@
The aforementioned extension allows for a new type of localisation-type results. Another type of localisation was studied in [7]. This approach is related to optimal transport of vector measures [9], [8] and extensions of Lipschitz maps [6, 13].
The notions of generalised convexity and of submartingale transport of measures, explored in [12], have found its applications in a study of the monopolist’s problem by Figalli, Kim and McCann [17] and by McCann and Zhang in [28].
1.9 Structure of the paper
In Section 2 we provide several definitions and lemmata. In Section 3 we define the Gelfand transform and study its basic properties. In Section 4 we provide definition of convexity with respect to a convex cone of functions. Proposition 4.13 shows a characterisation of closed convex hulls with respect to a linear space of real-parts of a complex algebra. In Section 5 we prove the main results, Theorem 1.11, Theorem 1.18 and Theorem 1.23. We also elaborate on the analogy of Theorem 1.11 and the Krein–Šmulian theorem. In Section 6 we show that the lemma of Bremermann and Lelong implies the resolution of the Levi problem.
2 Preliminaries
@@
We shall say that a topological space $\Omega$ is exhaustible by compacta if there exists a sequence $(K_{i})_{i=1}^{\infty}$ of compact subsets of $\Omega$ such that
@@
It is readily seen that the condition that $\Omega$ is exhaustible by compacta is equivalent to $\Omega$ being $\sigma$-compact and locally compact.
@@
The family of semi-norms $(\lVert\cdot\rVert_{\mathcal{C}(K)})_{K\in\mathcal{K}}$ on $\mathcal{A}$ defines a locally convex topology on $\mathcal{A}$, which we shall denote by $\tau_{\mathcal{K}}$.
@@
If $\Omega$ is $\sigma$-compact, then $\mathcal{A}$ in $\tau_{\mathcal{K}}$ is metrisable, as the topology is induced by a countable separating family of semi-norms [36, Remark 1.38, c), p. 28]. If moreover $\Omega$ is separable, then for each compact set $K\subset\Omega$ the space $\mathcal{C}(K)$ is separable, as follows by the Stone–Weierstrass theorem, so that $\mathcal{A}$ is separable. Therefore separability and $\sigma$-compactness of $\Omega$ imply that $\mathcal{A}^{*}$ in the weak* topology is metrisable on its compact subsets.
@@
Let $\Omega$ be a set. Let $\mathcal{A}$ be a set of functions on $\Omega$ with values in (,]($\infty$,$\infty$]( - ∞ , ∞ ]. A set $\mathcal{L}$ of functions on $\Omega$ is said to be:
@@
    a convex cone whenever $\alpha f+\beta g\in\mathcal{L}$ for any $f,g\in\mathcal{L}$ and any numbers $\alpha,\beta\geq 0$;
@@
    stable under suprema provided that the supremum $\sup\{f_{i}\mid i\in I\}$ of any family $(f_{i})_{i\in I}\subset\mathcal{L}$ belongs to $\mathcal{L}$;
@@
    separating points of $\Omega$ whenever for any two distinct $\omega_{1},\omega_{2}\in\Omega$, there exists $f\in\mathcal{L}$ such that $f(\omega_{1})\neq f(\omega_{2})$.
Let us stress that we allow the functions to take the value +$\infty$+ ∞. Note however that this is not allowed for functions in linear subspaces.
@@
$f(\omega)=\sup\{a_{i}(\omega)\mid i\in I\}\text{ for all }\omega\in\Omega.$ (2.1)
Proof 2.9.
@@
Clearly, $\mathcal{F}\supset\mathcal{G}$. Since $\mathcal{G}$ is a complete lattice cone, $\mathcal{F}=\mathcal{G}$.
@@
If $Y$ is the space of all linear, continuous functionals on a linear topological space $X$, then the complete lattice cone generated by $Y$ is the set of all convex, lower semi-continuous functions on $X$. We refer the reader to [32, Chapter 3, p. 13] for a proof of this fact.
@@
Let $\mathcal{F}$ be the complete lattice cone of functions on $\Omega$ that contains constants. Let $p\in\mathcal{F}$. Let $\xi\colon\mathbb{R}\to\mathbb{R}$ be convex and non-decreasing. Then the composition $\xi(p)$ belongs to $\mathcal{F}$.
Proof 2.12.
@@
For any $j\in J$, there exist $\alpha_{j}\geq 0$ and $\beta_{j}\in\mathbb{R}$ such that $\lambda_{j}(t)=\alpha_{j}t+\beta_{j}$ for all $t\in\mathbb{R}$, so that $\alpha_{j}p+\beta_{j}\in\mathcal{F}$ for each $j\in J$. Thus $\xi(p)\in\mathcal{F}$.
@@
Let $\Omega$ be a topological space. A map $p\colon\Omega\to\mathbb{R}$ is said to be proper if the preimage $p^{-1}(Z)$ ofr every compact sets in $Z\subset\mathbb{R}$ is compact in $\Omega$. Let $\mathcal{F}$ be a cone of functions on $\Omega$. We say that
@@
3 Gelfand transform
Given a topological space $\Omega$ and a linear space $\mathcal{A}$ of continuous functions on $\Omega$, one may consider the space of all linear functionals on $\mathcal{A}$ that are evaluations at points of $\Omega$. Let $\Phi\colon\Omega\to\mathcal{A}^{*}$ be given by the formula $\Phi(\omega)(a)=a(\omega)$ for $\omega\in\Omega$ and $a\in\mathcal{A}$. We shall call $\Phi$ the Gelfand transform, in analogy with the notion studied in the theory of Banach algebras, cf. [43].